OpenStudy (loser66):

Determine value of c for which, the function is a joint pdf of 2 continuous random variables X and Y f(x,y) = cxy \(0\leq x\leq 1\) \(x^2\leq y\leq x\) Please, help

OpenStudy (loser66):

To find c, I need \[\int_{0}^{1}\int_{x^2}^{x}cxy dydx\\ c\int_{0}^{1}\dfrac{xy^2}{2}dx |_{x^2}^{x} = \dfrac{c}{2}\int_{0}^{1}0dx\] Which is wrong. But I don't know why. The answer from the book is c = 24